A is two times as old as B. 8 years ago A was 3 times as old as B. how old is A and B now?
Answer by mananth(16949)

b's age x
A's age =2x
8 years ago
B's age = (x-8)
A's age = (2x-8)
(2x-8)=3(x-8)
2x-8 = 3x-24
x=16
B's age 16 years
A's age 32 years